The apparent magnitudes of two stars are 0.06 and 1.06 respectively. Calculate the ratio of their brightness.
1
Expert's answer
2019-05-29T09:31:16-0400

Pogson's formula is given by

"m_2-m_1=2.512\lg\frac{I_1}{I_2}"

So, ratio of stars brightness

"\frac{I_1}{I_2}=10^{0.4(m_2-m_1)}"

"=10^{0.4(1.06-0.06)}=2.512"